# Dynamical Low-Rank Compression of Neural Networks with Robustness under Adversarial Attacks

### v1

@@ 5 Numerical Results
[…]
@@
UCM dataset: We observe in Table 2 that Algorithm 1 can compress the VGG11, VGG16 and ViT-32b networks equally well as the non-regularized low-rank compression and achieves the first goal of high compression values of up to $95\%$ reduction of trainable parameters. Furthermore, the clean accuracy is similar to the non-compressed baseline architecture; thus, we achieve the second goal of (almost) loss-less compression. Noting the adversarial accuracy results under the $\ell^{2}$-FGSM, Jitter, and Mixup attacks with various attack strengths $\epsilon$, we observe that across all tests, the regularized low-rank network of Algorithm 1 significantly outperforms the non-regularized low-rank network. For the $\ell^{2}$-FGSM attack, our method is able to recover the adversarial accuracy of the baseline network. For Mixup, the regularization almost doubles the baseline accuracy for VGG16. By targeting the condition number of the weights, which gives a bound on the relative growth of the loss w.r.t. the size of the input, we postulate that the large improvement could be attributed to the improved robustness against the scale invariance attack (lin2020nesterov, , Section 3.3) included in MixupWe refer the reader to Section A.1.4 for a precise definition of the Mixup attack featuring scale invariance. However, this hypothesis was not further explored and is delayed to a future work. Finally, we are able to recover half of the lost accuracy in the Jitter attack. Overall, we achieved the third goal of significantly increasing adversarial robustness of the compressed networks. We refer to Table 6 for the used values of $\beta$ and to Section A.3 for extended numerical results for a grid of different $\beta$ values.
@@
Table 3: Comparison to literature on CIFAR10 with VGG16 under the $\ell^{1}$-FGSM. The first three rows list the computed mean over $10$ random initializations. The values of all other methods, given below the double rule, are taken from (NEURIPS2024_ea48cb23, , Table 1). RobustDLRT has higher adversarial accuracy at higher compression rates than all listed methods.
@@
CondLR savostianova2023robustlowranktrainingapproximate $\xi=0.5$ 50 89.97 72.25 60.19 50.17
CondLR savostianova2023robustlowranktrainingapproximate $\xi=0.5$ 80 89.33 68.23 48.54 36.66
@@
SVD prune yang2020learninglowrankdeepneural 50 89.92 67.30 47.77 36.98
SVD prune yang2020learninglowrankdeepneural 80 87.99 63.57 42.06 29.27
Cifar10 dataset: We repeat the methodology of the UCM dataset for Cifar10, and observe similar computational results in Table 2. Furthermore, we compare our method in Table 3 to several methods of the recent literature, see Section 2. We compare the adversarial accuracy under the $\ell^{1}$-FGSM attack, see Section A.1.2 for details, for consistency with the literature results. We find that our proposed method achieves the highest adversarial validation accuracy for all attack strengths $\epsilon$, even surpassing the baseline adversarial accuracy. Additionally, we find an at least 15% higher compression ratio with RobustDLRT than the second best compression method, CondLR savostianova2023robustlowranktrainingapproximate .
Table 4: UCM dataset – Black-box attack. Adversarial images with the $\ell^{2}$-FGSM attack are generated by the baseline network for different values of $\epsilon$. The baseline, low-rank, and regularized low-rank networks are evaluated on these images. Regularized low-rank compression achieves high adversarial accuracy, even under strong attacks.
@@
Black-box attacks: We investigate the scenario, where an attacker has knowledge of the used model architecture, but not of the low-rank compression. We use the Imagenet-1k pretrained VGG16 and VGG11 and re-train it with Algorithm 1 and baseline training on the UCM data using the same training hyperparameters. Then we generate adversarial examples with the baseline network and evaluate the performance on the low-rank network with and without regularization. The results are given in Table 4. In this scenario, the weights from low-rank training, being sufficiently far away from the baseline, provide an effective defense against the attack. Further, the proposed regularization significantly improves the adversarial robustness when compared to the unregularized low-rank network. Even for extreme attacks with $\epsilon=1$, the regularized network achieves $84.76\%$ and $87.33\%$ accuracy for VGG16 and VGG11 respectively.
@@ 6 Conclusion
RobustDLRT enables highly compressed neural networks with strong adversarial robustness by controlling the spectral properties of low-rank factors. The method is efficient and rank-adaptive that yields an over 94% parameter reduction across diverse models and attacks. The method achieves competitive accuracy, even for strong adversarial attacks, surpassing the current literature results by a significant margin. Therefore, we conclude the proposed method scores well in the combined metric of compression, accuracy and adversarial robustness.
@@
This research used resources of the Compute and Data Environment for Science (CADES) at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725"
Appendix A Details to the numerical experiments of this work
A.1 Recap of adversarial attacks
@@ A.1.1 $\ell^{2}$-FGSM attack
A.1.2 $\ell^{1}$-FGSM attack
The $\ell^{1}$-FGSM attack [37] is used in the reference work of [29] and uses the same workflow as (A.1.1), where (12) is changed to
@@ A.1.3 Jitter attack
A.1.4 Mixup attack
The Mixup attack is an adversarial attack that generates adversarial samples that share similar feature representations with an given virtual example [42]. Inspired by the Mixup data augmentation technique, this attack aims to create adversarial examples that maintain characteristics of both the original sample and its adversarial counterpart. Given a neural network classifier $f_{\theta}$ with parameters $\theta$, an input $x$, and its corresponding label $y$, the attack first computes a linear combination of cross-entropy and negative KL-divergence loss,
$\mathcal{L}_{\textup{mixup}}=\beta\sum_{k=1}^{5}\mathcal{L}_{\text{CE}}(f_{\theta}(\frac{x}{2^{k}}),y)-\mathcal{L}_{\text{KL}}$ (21)
@@ A.1.4 Mixup attack
Equation 21 features a scale invariance attack applied to the loss [22, Section 3.3].
@@
Table 5: Training hyperparameters for the UCM and Cifar10 Benchmark. The first set hyperparameters apply to both DLRT and baseline training, and we train DLRT with the same hyperparameters as the full-rank baseline models. The second set of hyper-parameters is specific to DLRT. The DLRT hyperparameters are selected by an initial parameter sweep. We choose the DLRT truncation tolerance relative to the Frobenius norm of ${\widehat{S}}$, i.e. $\vartheta=\tau\|{\widehat{S}}\|_{F}$, as suggested in [32].
Hyperparameter VGG16 VGG11 ViT
Batch Size (UCM) 16 16 16
Batch Size (Cifar10) 128 128 128
Learning Rate 0.001 0.001 0.001
Number of Epochs 20 20 5
L2 regularization 0 0 0.001
Optimizer Adam Adam Adam
DLRT rel. truncation tolerance $\tau$ 0.1 0.05 0.08
Coefficient Steps 10 10 10
Initial Rank 150 150 150
A.2 Network architecture and training details
@@
  • ViT32b is a Vision Transformer with 32x32 patch size, a deep learning architecture that leverages transformer models for image classification tasks.
The full training setup is described in Table 5. We train DLRT with the same hyperparameters as the full-rank baseline models. It is known [31] that DLRT methods are robust w.r.t. common hyperparameters as learning rate, and batch-size, and initial rank. The truncation tolerance $\tau$ is chosen between $0.05$ and $0.1$ per an initial parameter study. These values are good default values, as per recent literature [30, 35]. In general, there is a trade-off between target compression ratio and accuracy, as illustrated e.g. in [32] for matrix-valued and [35] for tensor-valued (CNN) layers.
A.3 UCM Test Case
Table 6: Overview of the $\beta$ for best performing regularization strength for RobustDLRTof Table 2.
UCM Dataset Cifar10 Dataset
Architecture FGSM Jitter Mixup FGSM Jitter Mixup
Vgg16 0.075 0.2 0.15 0.05 0.05 0.05
Vgg11 0.1 0.05 0.15 0.15 0.05 0.2
ViT-32b 0.1 0.15 0.15 0.01 0.01 0.05
The UC Merced (UCM) Land Use Dataset is a benchmark dataset in remote sensing and computer vision, introduced in [44]. It comprises 2,100 high-resolution aerial RGB images, each measuring 256×256 pixels, categorized into 21 land use classes with 100 images per class. The images were manually extracted from the USGS National Map Urban Area Imagery collection, covering various urban areas across the United States. The dataset contains images with spatial resolution approximately 0.3 meters per pixel (equivalent to 1 foot), providing detailed visual information suitable for fine-grained scene classification tasks.
@@ A.3 UCM Test Case
The numerical results for the whitebox FGSM, Jitter, and Mixup adversarial attacks on the VGG16 and VGG11 architectures can be found in Figure 4, Figure 5, and Figure 7. The regularizer confidently increases the adversarial validation accuracy of the networks.
In Table 7, we observe that the regularizer ${\mathcal{R}}(W)$ applied to the full weight matrices (and flattened tensors) $W$ in baseline format is able to increase the adversarial robustness of the baseline network in the UCM/VGG16 test case. However, the increased adversarial robustness comes at the expense of some of the clean validation accuracy.
@@
Figure 4: UCM Dataset, VGG16 clean and adversarial accuracy under the FGSM attack. Data is averaged over 10 stochastic training runs. The top row displays the full baseline network with $0\%$ c.r. and the matrix below displays the low-rank and regularized networks trained with Algorithm 1. All numbers display the mean of $10$ randomized training runs, where the randomness stems from shuffled batches. The initial condition of all runs is given by Imagenet-1k pretrained weights. The regularized low-rank networks with $\beta=0.075$ are able to recover the adversarial robustness of the baseline training while compressed by 95.84%. Results for VGG11 and Vit32b are similar.
Figure 5: UCM Dataset, VGG16 clean and adversarial accuracy under the Jitter attack. Data is averaged over 10 stochastic training runs. The top row displays the full baseline network with $0\%$ c.r. and the matrix below displays the low-rank and regularized networks trained with Algorithm 1. All numbers display the mean of $10$ randomized training runs, where the randomness stems from shuffled batches. The initial condition of all runs is given by Imagenet-1k pretrained weights. The regularized low-rank networks are able to recover most of the adversarial robustness of the baseline network. Results for VGG11 and Vit32b are similar.
Figure 6: VGG16
Figure 7: UCM Dataset, VGG16 clean and adversarial accuracy under the Mixup attack. Data is averaged over 10 stochastic training runs. The top row displays the full baseline network with $0\%$ c.r. and the matrix below displays the low-rank and regularized networks trained with Algorithm 1. All numbers display the mean of $10$ randomized training runs, where the randomness stems from shuffled batches. The initial condition of all runs is given by Imagenet-1k pretrained weights. The regularized low-rank networks almost double the adversarial accuracy of the baseline network at $95.84\%$ compression rate. Results for VGG11 and Vit32b are similar.
A.4 Cifar10

### v2

@@ 6 Numerical Results
We evaluate the numerical performance of Algorithm˜1 compared with non-regularized low-rank training, baseline training, and several other robustness-enhancing methods the VGG16, VGG11, and ViT-16b architectures and University of California, Merced (UCM), Cifar10, and ImageNet1k datasets. Detailed descriptions of the models, datasets, pre-processing, training hyperparameters, and competitor methods are given in Appendix˜B. We measure the compression rate (c.r.) as the relative amount of pruned parameters of the target network, i.e. $\textup{c.r.}=(1-\frac{\#\textup{params low-rank net}}{\#\textup{params baseline net}})\times 100$. The reported numbers in the tables represent the average over 10 stochastic training runs. We observe in Table˜2 that clean accuracy results exhibit a standard deviation of less than 0.8%; the standard deviation increases with the attack strength $\epsilon$ for all tests and methods. This observation holds true for all presented results; thus, we omit the error bars in the other tables for the sake of readability.
Table 3: Imagenet Benchmark, ViT-32l trained with baseline Adam, DLRT, and RobustDLRT. We report the low-rank results for unregularized $\beta=0.0$ and the best performing $\beta$, given in Table˜9. Algorithm˜1 (RobustDLRT) is able to match or surpass baseline adversarial accuracy values in most setups. All runs where RobustDLRT surpasses the uncompressed baseline are highlighted.
Top1/Top5 Clean Top1/Top5 Acc [%] for $\ell^{2}$-FGSM, $\epsilon$ Top1/Top5 Acc [%] for Jitter, $\epsilon$
Method c.r. [%] Acc. [%] 0.05 0.1 0.3 0.035 0.045
Baseline 0 74.37/92.20 43.58/73.75 31.42/63.42 16.03/43.41 43.09/78.24 35.57/74.96
DLRT 58.02 72.27/90.06 42.70/70.43 30.32/60.90 15.47/40.58 43.98/74.49 38.44/ 71.31
RobustDLRT 57.98 72.25/90.03 43.17/71.58 35.11/62.82 25.24/50.65 48.22/77.35 43.51/75.14
@@
UCM dataset We observe in Table˜2 that Algorithm˜1 can compress the VGG11, VGG16 and ViT-16b networks equally well as the non-regularized low-rank compression and achieves the first goal of high compression values of up to $95\%$ reduction of trainable parameters. Furthermore, the clean accuracy is similar to the non-compressed baseline architecture; thus, we achieve the second goal of (almost) loss-less compression. Noting the adversarial accuracy results under the $\ell^{2}$-FGSM, Jitter, and Mixup attacks with various attack strengths $\epsilon$, we observe that across all tests, the regularized low-rank network of Algorithm˜1 significantly outperforms the non-regularized low-rank network. For the $\ell^{2}$-FGSM attack, our method is able to recover the adversarial accuracy of the baseline network. For Mixup, the regularization almost doubles the baseline accuracy for VGG16. By targeting the condition number of the weights, which gives a bound on the relative growth of the loss w.r.t. the size of the input, we postulate that the large improvement could be attributed to the improved robustness against the scale invariance attack (lin2020nesterov, , Section 3.3) included in Mixup. We refer the reader to Section˜B.1.4 for a precise definition of the Mixup attack featuring scale invariance. However, this hypothesis was not further explored and is delayed to a future work. Finally, we are able to recover half of the lost accuracy in the Jitter attack. Overall, we achieved the third goal of significantly increasing adversarial robustness of the compressed networks. We refer to Table˜9 for the used values of $\beta$ and to Section˜B.3 for extended numerical results for a grid of different $\beta$ values.
@@
Table 4: Comparison to literature on CIFAR10 with VGG16 under the $\ell^{1}$-FGSM attack. The first three rows list the computed mean over $10$ random initializations. The values of all other methods, given below the double rule, are taken from (savostianova2023robustlowranktrainingapproximate, , Table 1). RobustDLRT has higher adversarial accuracy at higher compression rates than all listed methods.
@@
CondLR savostianova2023robustlowranktrainingapproximate $\tau=0.5$ 50 89.97 72.25 60.19 50.17
CondLR savostianova2023robustlowranktrainingapproximate $\tau=0.5$ 80 89.33 68.23 48.54 36.66
@@
SVD prune yang2020learning 50 89.92 67.30 47.77 36.98
SVD prune yang2020learning 80 87.99 63.57 42.06 29.27
Cifar10 dataset We repeat the methodology of the UCM dataset for Cifar10, and observe similar computational results in Table˜2. Furthermore, we compare our method in Table˜4 to several methods of the recent literature, see Section˜3. We compare the adversarial accuracy under the $\ell^{1}$-FGSM attack, see Section˜B.1.2 for details, for consistency with the literature results. We find that our proposed method achieves the highest adversarial validation accuracy for all attack strengths $\epsilon$, even surpassing the baseline adversarial accuracy. Additionally, we find an at least 15% higher compression ratio with RobustDLRT than the second best compression method, CondLR savostianova2023robustlowranktrainingapproximate . A similar experiment for the Projected Gradient Descent (PGD) attack madry2018towards is given in Appendix˜A.
ImageNet1k dataset Finally we repeat the methodology for the ImageNet1k dataset using the ViT-32l vision transformer trained from an ImageNet21k checkpoint, and report the results in Table˜3. The hyperparameters are obtained by an initial sweep and reported in Tables˜8 and 9. RobustDLRT consistently yields higher Top-1/Top-5 accuracy across FGSM and Jitter attacks than DLRT, with especially pronounced gains at larger perturbations (e.g., +9+9 points in Top-1 accuracy under FGSM $\epsilon=0.3$). These trends are consistent with our ViT experiments in Table˜2, demonstrating that adversarial regularization enhances robustness without compromising scalability. We benchmark the training runtime of one ImageNet epoch on an A100 80GB GPU. DLRT requires 26m 07s, while RobustDLRT (with the regularizer) requires 27m 51s, corresponding to an overhead of approximately 3%. This overhead can likely be reduced with further implementation optimizations, indicating that our approach is computationally scalable.
Black-box attacks We investigate the scenario where an attacker has knowledge of the used model architecture, but not of the low-rank compression. We use the Imagenet-1k pretrained VGG16 and VGG11 and re-train it with Algorithm˜1 and baseline training on the UCM data using the same training hyperparameters. Then we generate adversarial examples with the baseline network and evaluate the performance on the low-rank network with and without regularization. The results are given in Table˜5. In this scenario, the weights from low-rank training, being sufficiently far away from the baseline, provide an effective defense against the attack. Further, the proposed regularization significantly improves the adversarial robustness when compared to the unregularized low-rank network. Even for extreme attacks with $\epsilon=1$, the regularized network achieves $84.76\%$ and $87.33\%$ accuracy for VGG16 and VGG11 respectively.
Adversarial Training We evaluate the performance of low-rank training for VGG16 on the UCM dataset using adversarial training. Following goodfellow2014explaining , we use the FGSM attack for different values of $\epsilon$ and train on both 50% clean and attacked images per batch. The results reported in Table˜6 illustrate that RobustDLRT is both compatible with and able to benefit from adversarial training. DLRT without regularization benefits slightly from adversarial training, but exhibits a clear margin to RobustDLRT. RobustDLRT is able to approximately match the non-compressed baseline.
Table 5: UCM dataset – Black-box attack. Adversarial images with the $\ell^{2}$-FGSM attack are generated by the baseline network for different values of $\epsilon$. The baseline, DLRT $(\beta=0)$, and RobustDLRT $(\beta=0.075)$ networks are then evaluated on these images. Regularized low-rank compression achieves high adversarial accuracy, even under strong attacks.
@@
Table 6: UCM dataset – Adversarial Training. VGG16 is trained on 50% clean images and 50% images attacked with $\ell^{2}$-FGSM for various $\epsilon$. The displayed numbers are the mean of 5 repeated runs. RobustDLRT $(\beta=0.075)$ is superior to DLRT $(\beta=0)$ and is able to approximately match the non-compressed baseline.
$\ell^{2}$-FGSM, $\epsilon$
Method c.r. [%] 0.0 0.1 0.5 0.75 1.0
Baseline 0.0 92.61 91.91 91.90 89.61 89.91
$\beta=0$ 94.46 92.55 91.91 87.98 85.37 82.96
$\beta=0.075$ 94.19 92.49 92.49 90.98 89.56 89.42
@@ 7 Conclusion
RobustDLRT enables highly compressed neural networks with strong adversarial robustness by controlling the spectral properties of low-rank factors. The method is efficient, rank-adaptive, and yields an up to 94% parameter reduction across a diverse suite of models and datasets. The method achieves competitive accuracy, even for strong adversarial attacks, surpassing the current literature results by a significant margin. Therefore, we conclude the proposed method scores well in the combined metric of compression, accuracy and adversarial robustness.
@@
This research used resources of the Compute and Data Environment for Science (CADES) at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.
Appendix A Additional Numerical Results
UCM Dataset
The numerical results for the whitebox FGSM, Jitter, and Mixup adversarial attacks on the VGG16 and VGG11 architectures can be found in Figure˜4, Figure˜5, and Figure˜6. The regularizer confidently increases the adversarial validation accuracy of the networks.
In Table˜10, we observe that the regularizer ${\mathcal{R}}(W)$ applied to the full weight matrices (and flattened tensors) $W$ in baseline format is able to increase the adversarial robustness of the baseline network in the UCM/VGG16 test case. However, the increased adversarial robustness comes at the expense of some of the clean validation accuracy.
Cifar10 Dataset
We run the same experiment in Table˜4 but with the $\ell^{2}$-PGD attack, which is an iterative version of $\ell^{2}$-FGSM with an random perturbation of the input image as the initial condition. Overall, we see that RobustDLRT is competitive with the other robustness-improving methods when the compression rate is taken into account.
Table 7: Comparison to literature on CIFAR10 with VGG16 under the $\ell^{2}$-PGD attack. The first three rows list the computed mean over $10$ random initializations. The values of all other methods, given below the double rule, are taken from [35, Table 5]. RobustDLRT has competitive adversarial accuracy to all methods with a compression rate $\geq 80\%$.
$\ell^{2}$-PGD, $\epsilon$
Method c.r. [%] 0.0 0.1 0.13 0.16 0.2 0.23 0.26 0.3
RobustDLRT $\beta=0.15$ 94.18 88.80 62.58 53.47 44.95 34.75 28.33 22.64 16.59
DLRT 94.53 88.58 59.34 50.06 41.50 31.82 25.67 20.48 15.04
Baseline 0 90.48 63.01 54.66 47.87 40.77 36.75 33.51 29.93
Cayley SGD [25] 0 89.62 67.68 59.38 51.09 40.87 34.46 29.21 23.62
Projected SGD [1] 0 89.70 67.64 59.25 51.06 40.86 34.51 29.19 23.64
CondLR [35] $\tau=0.1$ 50 90.93 67.03 62.08 59.15 56.92 55.96 55.28 54.58
CondLR [35] $\tau=0.5$ 50 89.97 64.84 60.25 57.75 56.03 55.21 54.75 54.25
CondLR [35] $\tau=0.1$ 80 90.48 61.00 50.84 42.19 33.70 29.44 26.55 23.97
CondLR [35] $\tau=0.5$ 80 89.33 57.45 46.35 37.20 28.30 23.82 20.65 17.84
LoRA [17] 50 89.97 55.74 45.11 36.86 29.62 26.28 24.02 21.84
LoRA [17] 80 88.10 51.40 39.70 30.12 20.97 16.29 13.15 10.37
SVD prune [50] 50 89.92 54.87 43.85 35.23 27.95 24.38 22.06 19.94
SVD prune [50] 80 87.99 50.64 39.06 29.57 20.16 15.49 12.22 9.57
Appendix B Details to the numerical experiments of this work
B.1 Recap of adversarial attacks
@@ B.1.1 $\ell^{2}$-FGSM attack
B.1.2 $\ell^{1}$-FGSM attack
The $\ell^{1}$-FGSM attack [44] is used in the reference work of [35] and uses the same workflow as (B.1.1), where (12) is changed to
@@ B.1.3 Jitter attack
B.1.4 Mixup attack
The Mixup attack [49] is an adversarial attack that generates adversarial samples that share similar feature representations with an given virtual example. Inspired by the Mixup data augmentation technique, this attack aims to create adversarial examples that maintain characteristics of both the original sample and its adversarial counterpart. Given a neural network classifier $f_{\theta}$ with parameters $\theta$, an input $x$, and its corresponding label $y$, the attack first computes a linear combination of cross-entropy and negative KL-divergence loss,
$\mathcal{L}_{\textup{mixup}}=\beta\sum_{k=1}^{5}\mathcal{L}_{\text{CE}}\bigg(f_{\theta}\Big(\frac{x}{2^{k}}\Big),y\bigg)-\mathcal{L}_{\text{KL}}$ (21)
@@ B.1.4 Mixup attack
Equation˜21 features a scale invariance attack applied to the loss [27, Section 3.3].
@@
Table 8: Training hyperparameters for the UCM, Cifar10, and ImageNet Benchmarks. The first set hyperparameters apply to both DLRT and baseline training, and we train DLRT with the same hyperparameters as the full-rank baseline models. The second set of hyper-parameters is specific to DLRT. The DLRT hyperparameters are selected by an initial parameter sweep. We choose the DLRT truncation tolerance relative to the Frobenius norm of $^{\widehat{S}}$, i.e. $\vartheta=\tau\|{\widehat{S}}\|_{F}$, as suggested in [38].
Hyperparameter VGG16 VGG11 ViT16b ViT32l
Batch Size (UCM) 16 16 16 n.a.
Batch Size (Cifar10) 128 128 128 n.a.
Batch Size (ImageNet) n.a. n.a. n.a. 256
Learning Rate 0.001 0.001 0.001 0.001
Number of Epochs 20 20 5 10
L2 regularization 0 0 0.001 0.0001
Optimizer AdamW AdamW AdamW AdamW
DLRT rel. truncation tolerance $\tau$ 0.1 0.05 0.08 0.013
Coefficient Steps $s_{*}$ 10 10 10 75
Initial Rank 150 150 150 200
Parameters 138M 132M 86M 304M
B.2 Network architecture and training details
@@
  • ViT16b is a Vision Transformer with 16x16 patch size, a deep learning architecture that leverages transformer models for image classification tasks.
  • ViT32l is a Vision Transformer with 32x32 patch size, a deep learning architecture that leverages transformer models for image classification tasks. We use the Imagenet21k weights from the huggingface endpoint google/vit-large-patch32-224-in21k as weight initialization.
The full training setup is described in Table˜8. We train DLRT with the same hyperparameters as the full-rank baseline models. It is known [37] that DLRT methods are robust w.r.t. common hyperparameters as learning rate, and batch-size, and initial rank. The truncation tolerance $\tau$ is chosen between $0.05$ and $0.1$ per an initial parameter study. These values are good default values, as per recent literature [36, 42]. In general, there is a trade-off between target compression ratio and accuracy, as illustrated e.g. in [38] for matrix-valued and [42] for tensor-valued (CNN) layers.
B.3 UCM Test Case
Table 9: Overview of the $\beta$ for best performing regularization strength for RobustDLRT of Table˜2.
UCM Dataset Cifar10 Dataset ImageNet Dataset
Architecture FGSM Jitter Mixup FGSM Jitter Mixup FGSM Jitter Mixup
VGG16 0.075 0.2 0.15 0.05 0.05 0.05 n.a. n.a. n.a.
VGG11 0.1 0.05 0.15 0.15 0.05 0.2 n.a. n.a. n.a.
ViT16b 0.1 0.15 0.15 0.01 0.01 0.05 n.a. n.a. n.a.
ViT32l n.a. n.a. n.a. n.a. n.a. n.a. 0.075 0.075 0.075
The University of California, Merced (UCM) Land Use Dataset is a benchmark dataset in remote sensing and computer vision, introduced in [51]. It comprises 2,100 high-resolution aerial RGB images, each measuring 256×256 pixels, categorized into 21 land use classes with 100 images per class. The images were manually extracted from the USGS National Map Urban Area Imagery collection, covering various urban areas across the United States. The dataset contains images with spatial resolution approximately 0.3 meters per pixel (equivalent to 1 foot), providing detailed visual information suitable for fine-grained scene classification tasks.
@@ B.3 UCM Test Case
Figure 4: UCM Dataset, VGG16 clean and adversarial accuracy under the FGSM attack. Data is averaged over 10 stochastic training runs. The top row displays the full baseline network with $0\%$ c.r. and the matrix below displays the low-rank and regularized networks trained with Algorithm˜1. All numbers display the mean of $10$ randomized training runs, where the randomness stems from shuffled batches. The initial condition of all runs is given by Imagenet-1k pretrained weights. The regularized low-rank networks with $\beta=0.075$ are able to recover the adversarial robustness of the baseline training while compressed by 95.84%. Results for VGG11 and Vit16b are similar.
Figure 5: UCM Dataset, VGG16 clean and adversarial accuracy under the Jitter attack. Data is averaged over 10 stochastic training runs. The top row displays the full baseline network with $0\%$ c.r. and the matrix below displays the low-rank and regularized networks trained with Algorithm˜1. All numbers display the mean of $10$ randomized training runs, where the randomness stems from shuffled batches. The initial condition of all runs is given by Imagenet-1k pretrained weights. The regularized low-rank networks are able to recover most of the adversarial robustness of the baseline network. Results for VGG11 and Vit16b are similar.
Figure 6: UCM Dataset, VGG16 clean and adversarial accuracy under the Mixup attack. Data is averaged over 10 stochastic training runs. The top row displays the full baseline network with $0\%$ c.r. and the matrix below displays the low-rank and regularized networks trained with Algorithm˜1. All numbers display the mean of $10$ randomized training runs, where the randomness stems from shuffled batches. The initial condition of all runs is given by Imagenet-1k pretrained weights. The regularized low-rank networks almost double the adversarial accuracy of the baseline network at $95.84\%$ compression rate. Results for VGG11 and Vit16b are similar.
B.4 Cifar10